i need help.
write a system of equations.
The receipts from 300 tickets for a high school program were $650. Adults were charged $2.50 each and the students $1.50. How many tickets of each kind were sold?

Answer by rfer(16322)

a+s=300 s=300-a
---------------
2.50a+1.50s=650
2.50a+1.50(300-a)=650
2.50a+450-1.50a=650
a=200
-----------------
s=300-200
s=100